Resistors for electronic circuits are manufactured on a high-speed automated machine. The machine is set up to produce a large run of resistors of 1,000 ohms each. Use Exhibit 10.13.

To set up the machine and to create a control chart to be used throughout the run, 15 samples were taken with four resistors in each sample. The complete list of samples and their measured values are as follows: Use three-sigma control limits.

SAMPLE NUMBER READINGS (IN OHMS)
1 992 992 1000 1027
2 978 994 985 1018
3 1021 998 1018 994
4 974 1022 999 1026
5 1000 1012 1030 991
6 974 987 985 987
7 1005 973 1011 1001
8 983 1024 1012 977
9 990 1004 973 999
10 990 1018 1024 1001
11 998 976 1029 980
12 1024 1020 1020 976
13 1026 1018 1029 971
14 982 979 982 989
15 991 988 980 994

a. Calculate the mean and range for the above samples. (Round "Mean" to 2 decimal places and "Range" to the nearest whole number.)

Sample Number Mean Range
1
2
3
4
5
6
7
8
9
10
11
12
13
14
15

b. Determine X=X= and R−R−. (Round your answers to 3 decimal places.)

X=X=
R−R−

c. Determine the UCL and LCL for a X−X−chart. (Round your answers to 3 decimal places.)

UCL
LCL

d. Determine the UCL and LCL for R-chart. (Leave no cells blank - be certain to enter "0" wherever required. Round your answers to 3 decimal places.)

UCL
LCL

e. What comments can you make about the process?

The process is in statistical control.
The process is out of statistical control.
Answer #1

Answers a are in BOLD

sample 1 2 3 4 R= Max- Min X-bar= average of sample
1 992 992 1000 1027 35 1002.75
2 978 994 985 1018 40 993.75
3 1021 998 1018 994 27 1007.75
4 974 1022 999 1026 52 1005.25
5 1000 1012 1030 991 39 1008.25
6 974 987 985 987 13 983.25
7 1005 973 1011 1001 38 997.50
8 983 1024 1012 977 47 999.00
9 990 1004 973 999 31 991.50
10 990 1018 1024 1001 34 1008.25
11 998 976 1029 980 53 995.75
12 1024 1020 1020 976 48 1010.00
13 1026 1018 1029 971 58 1011.00
14 982 979 982 989 10 983.00
15 991 988 980 994 14 988.25

Answer b

35.933 999.017

AVERAGE(all X-bar samples)

R-bar X-bar-bar

Answer c:

control limits for x-bar, z=3
UCL= Xbar-bar+ z*STD 1027.725
CL= X-bar-bar 999.017
LCL= Xbarbar- z*STD 970.308
Standard deviation= STD= STDEV(select all X-bar values) 9.5695

Answer d:

R-bar
control limits for Range, z=3
UCL=R-bar+ z*STD 80.740
CL= R-bar 35.933
LCL=R-bar- z*STD 0 if negative -8.873 0.000
Standerd deviation= STD= STDEV(select all range values) 14.9354

Answer e:

The process is in control: As we can see that NO sample is beyond control limits in charts. Every sample is within the UCL and LCL